5. Two vehicles are approaching an intersection. One is a 2500-kg pickup traveling at 14.0m/s from east to west (the βˆ’π‘₯-direction), and the other is a 1500-kg sedan going from south to north (the +𝑦-direction) at 23.0 m/s. (a) Find the π‘₯- and 𝑦-components of the net momentum of this system. (b) What are the magnitude and direction of the net momentum?


1
Expert's answer
2021-09-29T09:55:41-0400

(a) "p_x=-2500\cdot14=35000\ (kg\cdot m/s)"


"p_y=-1500\cdot23=34500\ (kg\cdot m/s)"


(b) "p=\sqrt{35000^2+34500^2}=49145\ (kg\cdot m/s)"


"\tan\alpha=\frac{p_y}{p_x}=\frac{34500}{-35000}=-0.9857\to \alpha=135.4°" (relative to the positive direction of the x-axis)